OpenStudy (adamk):

The half-life of carbon-14 is approximately 6000 years. Scientists obtain 100kg. Find a formula to represent this equation.

OpenStudy (adamk):

It's in the form f(x) = ab^x. I think you use the two points (0,100) and (6000, 1).

OpenStudy (perl):

actually in 6000 years it will be half of 100 , or 50 kg

OpenStudy (perl):

so (0,100) and (6000, 50)

OpenStudy (adamk):

Ah, OK. Thanks.

OpenStudy (adamk):

Still stuck on the rest.

OpenStudy (perl):

did you try to plug in values?

OpenStudy (adamk):

Yes. I did 50 = ab^6000. Then, a = 50/b^6000. Then, 100 = 50/b^6000(b)^0 100 = 50/b^6000 100b^6000 = 50 b^6000 = 1/2

OpenStudy (perl):

carbon-14 is approximately 6000 years. Scientists obtain 100kg. Find a formula to represent this equation. so we want to use y = a*b^x we have points (0,100 ) and (6000 , 50) first plug in (0,100) in y = a*b^x 100 = a * b^0 100 = a*1 100 = a This gives us y = 100 * b^x plug in (6000 , 50) 50= 100 * b^(6000) 50/100 = b^(6000) The left side simplifies to 1/2 or 0.5 0.5 = b^(6000) take the power of 1/6000 to both sides (0.5) ^(1/6000) = (b^6000)^(1/6000) by laws of exponents the right side cancels (0.5) ^(1/6000) = b^(6000*1/6000) (0.5) ^(1/6000) = b^1 (0.5) ^(1/6000) = b Therefore we have y = 100 * (0.5 ^(1/6000)) ^ x Now you have to options, you can evaluate 0.5 ^ (1/6000) which is approximately 0.99988 or you can leave it and multiply by the exponent x. option 1: y = 100 * (0.99988)^x option 2: y = 100 * (0.5 ^(1/6000)) ^ x or y = 100 * 0.5 ^ (x/6000)